OpenStudy (princessaurora):

** Will Fan and Medal ** What is the asymptote of the function: f(x) = (3^x + 1) – 2? A) y=-2 B) y=-1 C) y=1 D) y=2

OpenStudy (anonymous):

Y=-1

OpenStudy (princessaurora):

How did you get that answer? Please explain

OpenStudy (princessaurora):

@whpalmer4 is this right?

OpenStudy (whpalmer4):

Yes, \(y=-1\) is correct. Think about what happens to the value of \(f(x)\) as \(x\) moves to the left of \(0\): \[f(-1) = 3^{-1}+1-2 = \frac{1}{3}-1\] \[f(-2) = 3^{-2} +1 -2 = \frac{1}{3^2} -1 \] \[f(-100) = 3^{-100}+1-2 = \frac{1}{3^{100}}-1\]That \(3^x\) term gets arbitrarily close to \(0\), so the overall value gets arbitrarily close to \(y = f(x) = -1\)

OpenStudy (whpalmer4):

@PrincessAurora

OpenStudy (princessaurora):

@whpalmer4 thanks a lot for the help :)

OpenStudy (anonymous):

i gave u the answer

OpenStudy (anonymous):

@PrincessAurora

OpenStudy (whpalmer4):

@lildanny2025 OpenStudy isn't a place where you go to "get answers", it's a place where you go to learn how to find your own answers. The OpenStudy code of conduct ( http://openstudy.com/code-of-conduct) includes the line: "Don’t post only answers - guide the asker to a solution." @PrincessAurora specifically asked how you got that answer, but you didn't provide that information, which is the only valuable part of the exercise. How is knowing the answer to this particular problem is \(y=-1\) (without knowing how to find it) going to be any help solving the next problem of this sort?